A research company surveyed 1022 adults from one country (assume a 50/50 split between men and women) and found that 66% of men and 41% of women support scrapping the penny. Let men from this country be population 1 and women from this country be population 2. Complete parts a) through c) below. a) Construct a 95% confidence interval for the difference in support for scrapping the penny between men and women. Select the correct choice below and, if necessary, fill in the answer boxes within your choice. A. The 95% confidence interval is between | 18.1 % and 29.91% (Round to one decimal place as needed. Use ascending order.) O B. The interval should not be calculated because the assumptions and conditions are not met and cannot be reasonably assumed to be met.
Answer #1

Answer

(A) Formula for the confidence interval between proportion difference is given as

CI = (p_1-p_2) pm z*sqrt{[(p_1*(1-p_1))/n_1]+[(p_2*(1-p_2))/n_2]}

where p1 = 66% = 66/100= 0.66

p2= 41% = 41/100 = 0.41

sample size n1 = n2 = 1022/2 = 511

z score for 95% confidence interval is 1.96 (using z distribution table)

setting all the given values, we get

CI (0.66-0.41)±1.96*V/[0.66 * (1-0.66))/511| + [0.41 * (1-0.41))/5111

this gives us

C1 = (0.25)士1.96 * V0.0009 125 = 0.25 ± 0.0592

on solving, we get

CI = (0.191,0.309)

converting it into percentages, we get

0.191*100 =19.1% and 0.309*100=30.9%

So, required confidence interval is 19.1% and 30.9%